# Radial migration in the Galactic disc driven by a slowing bar

### v1

@@ 4.1 Age-metallicity sequences in the solar vicinity
Panel (a) of Fig. 1 shows the column-normalised age-metallicity distribution of the simulated particles in the solar vicinity. Two clear sequences are apparent: an upper sequence that rises rapidly at early times and flattens for stars formed since the bar began decelerating (marked by the orange-solid line), and a lower sequence that exhibits a strong age-metallicity correlation all the way to young ages. We identify the two sequences with migrated and local stars respectively. The red solid line, $\mathrm{[Fe/H]}(t,R_{\odot})$, represents the evolution of metallicity for particles that formed at the present solar radius, $R_{\odot}$, demonstrating that the local sequence is composed of stars born around the solar circle.
@@
To summarise, the migrated sequence is composed of the age-metallicity relation at the initial corotation radius for the time before the bar formation and the chunks of age-metallicity relations at the corresponding corotation radii at later times. The migrated sequence after the bar formation is flatter than the expected age-metallicity sequence at a fixed radius because stars at larger radii are dragged, but the extent of flattening depends on the details of the chemical evolution history (see Fig. A1 and Section 4.3.1 for clarification).
@@
In panel (c) of Fig. 1, we show the column-normalised age-metallicity distribution for the subgiant star sample described in Section 2. As for the test particle simulation, there are two distinct sequences marked by the blue and orange dashed lines. The test particle simulation results suggest that the lower sequence is composed of stars that formed around $R\sim R_{\odot}$, and the upper sequence is composed of stars formed in the inner Galaxy that have migrated with the expansion of the corotation radius. Xiang & Rix (2022) also pointed out the migrated origin of the upper sequence. Under this scenario, the turning point in the upper sequence, indicated by the red dashed line, gives an estimate of the bar formation epoch of $\sim 8$ Gyr ago, which is broadly consistent with other measurements of the bar formation time (Sanders et al., 2024; Haywood et al., 2024). We indicate the metallicity and age of the Sun with the red circle in panel (c). The Sun resides on the migrated sequence, suggesting that the Sun may have migrated to its current location with the corotation of the bar (Baba et al., 2024). A difference between panels (a) and (c) is that the upper envelope of the migrated sequence in the observations is decreasing slightly whilst its counterpart in the test particle simulation is flat. This implies that either the deceleration rate is underestimated (see discussion in Section 4.3.1) or other non-axisymmetric perturbations neglected in the simulation have enhanced the radial migration.
@@ 4.2 Implication for the disc $[\alpha/\mathrm{Fe}]$-bimodality
Our results supplement previous analyses that have shown that radial migration can reproduce the $[\alpha/\mathrm{Fe}]$-bimodality in the Galactic disc (Sharma et al., 2021a; Chen et al., 2023). The scale length growth of the disc has to be more finely tuned to match the observed radial dependence of the $\mathrm{[\alpha/Fe]-[Fe/H]}$ plane for radial migration mechanisms that do not favour the current corotation radius. Our analysis suggests decelerating bar-driven radial migration is the key mechanism responsible for the disc $[\alpha/\mathrm{Fe}]$-bimodality, which causes the migrated and local stars to co-exist around the present corotation radius, naturally. However, other effects in addition to radial migration could also cause $[\alpha/\mathrm{Fe}]$-bimodality in the Galactic disc, and our results cannot rule out their existence (e.g. Khoperskov et al., 2021).
@@ 5.1 Predictions for the bar-driven radial migration scenario
Azimuthal metallicity variation: Di Matteo et al. (2013) showed that ongoing radial migration can cause azimuthal abundance variations in the galactic disc. A key difference in the bar-driven radial migration scenario from other migration mechanisms is the azimuthal distribution of the migrated stars. Due to the characteristic shape of the corotation orbits, corotating stars naturally clump along the bar minor axis around the corotation radius. Therefore, the stars that experienced corotation dragging and migrated from the inner Galaxy would form an overdensity along the bar minor axis, causing the mean metallicity to peak along the bar minor axis. Azimuthal metallicity variation has been observed in the Milky Way disc (Poggio et al., 2022; Hawkins, 2023; Hackshaw et al., 2024). Recently, Khoperskov et al. (2024) measured the metallicity distribution of the entire Galactic disc with orbital reconstruction. As shown in fig. 22 therein, the metallicity for stars around guiding radii of $7-8$ kpc varies sinusoidally in the azimuthal direction and peaks along the bar minor axes. Such an azimuthal trend is not seen in stars at other guiding radii. This result is in agreement with the prediction of our model. However, kinematic fragmentation (Fragkoudi et al., 2017; Debattista et al., 2017; Zhang et al., 2024c) could also cause a similar result, as kinematically cold (more metal-rich) stars are easier to be trapped. Therefore, more analysis is needed to differentiate these mechanisms.
@@
Azimuthal variation in the high-$\alpha$ star fraction: As we discussed in Section 4.2, high-$\alpha$ stars observed in the solar vicinity are members of the migrated star sequence. Therefore, similarly to the azimuthal variation of metallicity, we might also expect an azimuthal variation for the high-$\alpha$ star fraction ($N_{\rm high-\alpha}$/$N_{\rm total}$), where more high-$\alpha$ stars are expected along the bar minor axes. We examine this using the sample we adopted in this work and a sample from APOGEE DR17. We indeed see an azimuthal variation in the fraction of high-$\alpha$ stars around the corotation radius (see Fig. D1). THe fraction is smallest around the bar’s major axis and increases towards the minor axis. The prediction is also in agreement with the results shown in Khoperskov et al. (2024), who found that the high-$\alpha$ sequence contributes more in the direction of the minor axes than the major axes.
@@
Coldness of the Galactic disc: Frankel et al. (2020) found that diffusion in the angular momentum of the Galactic disc stars is much greater than the scattering of radial actions. Recently, Hamilton et al. (2024) showed that it is difficult to transport stars with such coldness with transient spiral arms unless the lifetime and pitch angle for the spiral arms are fine-tuned (but also see the counterevidence in Sellwood & Binney, 2025). However, in the bar-driven radial migration scenario, the angular momentum of the corotating stars increases with the slowing of the bar, but the radial actions are invariant (see Sellwood & Binney 2002; Chiba et al. 2021 for theoretical explanations). Computationally, Khoperskov et al. (2020) showed that the orbits of stars transported with the corotation resonance can remain circular. So, the bar-driven radial migration scenario could alleviate the problem of the coldness of the migration, but more experiments or quantitative analysis are required to demonstrate this.
@@
Density profile of the disc: As argued, the Galactic high-$\alpha$ and high-$\mathrm{Fe}$ disc should also experience significant radial migration driven by the decelerating bar. Stars that experienced resonant dragging clump around the present-day corotation radius producing a maximum in the radial exponential surface density profile of the high-$\alpha$ and high-$\mathrm{Fe}$ disc leading to the broken-power law profiles as observed by (Yu et al., 2021; Lian et al., 2022b). More studies are required to quantify the correlation between disc density profiles and the bar-driven radial migration mechanism.
@@ 5.2 Caveats of the test particle simulation
Secondly, we also adopted a specific chemical model to assign the metallicity of the particles in the simulation. The detailed quantitative results presented in this work depend on the method in Lu et al. (2024b). However, the nature of our argument is that the decelerating bar-driven radial migration causes stars from very different birth radii to co-exist around the corotation radius, and the metallicity of the migrated sequence naturally flattens after the bar formation because the corotation resonance moves outward. Hence, switching to a different chemical model does not affect our argument as long as the radial metallicity gradient in the model is reasonable. Furthermore, our motivational calculation in Section 4.3.1 and the observation of the break in the upper migrated age-metallicity sequence are direct empirical indicators which lead to very similar conclusions as the more detailed birth radii calculation.
@@ 6 Conclusion
We provided a series of other predictions if the slowing down of the Galactic bar is the main driver responsible for the radial migration in the Milky Way disc. We expect that: (1) the mean metallicity of stars is higher along the bar’s minor axis around the corotation radius; (2) the fraction of high-$\alpha$ stars is the smallest along the bar’s major axis and increases towards the minor axis around the corotation radius, which is demonstrated in Fig. D1 using APOGEE survey; (3) the migrated stars experienced much more radial migration than radial heating; (4) the radial density profile of the high-$\alpha$ and high-Fe stars should be consistent with a broken power law with the break close to the corotation radius. As discussed in Section 5.1, many of these predictions are consistent with observations in the Milky Way.
@@ A.1 Time evolution of the galactic potential
where the initial pattern speed is $\Omega_{i}=80$ km s-1 kpc-1, final pattern speed $\Omega_{i}=34$ km s-1 kpc-1, $t_{0}\sim 4$ Gyr, $t_{1}\sim 5$ Gyr, $t_{2}\sim 5.5$ Gyr, and $t_{4}\sim 11.2$ Gyr. We also adjust the length of the bar scaled to its corotation radius so that the length of the bar matches its original value when the pattern speed is $39\,\mathrm{km\,s^{-1}\,kpc^{-1}}$ (the pattern speed value of the Portail et al. (2017) model). At the end of the simulation, the pattern speed is $34\,\mathrm{km\,s^{-1}\,kpc^{-1}}$.
@@ Appendix D Azimuthal variation for high-$\alpha$ star fraction
Figure D1: The azimuthal variation of the high-$\alpha$ star fraction. The blue, orange and green lines is for the guiding radii bins of $43<R_{g}/{\rm kpc}<$ 4, $76<R_{g}/{\rm kpc}<$ 7, $87<R_{g}/{\rm kpc}<$ 8, respectively. The grey dashed verticle line denotes the orientation of the bar major axis. A constant offset is added to each line for a better visualisation.
@@
We classify stars as a high-$\alpha$ star if $[\alpha/\rm M]>0.18$ for $[{\rm M/H}]<-0.7$ and $[\alpha/\rm M]>0.18-0.01([{\rm M/H}]+0.7)$ for $[{\rm M/H}]\geq-0.7$. We then take the ratio of the high-$\alpha$ to the total stars at different azimuthal angles in each of the guiding radii bins as shown in Fig. D1. There is an obvious increase of the high-$\alpha$ star fraction for the azimuthal angles away from the bar-major axis around the present-day guiding radii ($76<R_{g}/{\rm kpc}<$ 7), while the variation is much more insignificant inside and outside the corotation. This is in agreement with the expected consequence of the radial migration driven by the decelerating bar (see more discussion in Section 5.1).

### v2

# Observational constraints of radial migration in the Galactic disc driven by the slowing bar

@@ 4.1 Age-metallicity sequences in the solar vicinity
Panel (a) of Fig. 1 shows the column-normalised age-metallicity distribution of the simulated particles in the solar vicinity. Two clear sequences are apparent: an upper sequence that rises rapidly at early times and flattens for stars formed since the bar began decelerating (marked by the red-solid line), and a lower sequence that exhibits a strong age-metallicity correlation all the way to young ages. We identify the two sequences with migrated and local stars respectively. The orange solid line, $\mathrm{[Fe/H]}(t,R_{\odot})$, represents the evolution of metallicity for particles that formed at the present solar radius, $R_{\odot}$, demonstrating that the local sequence is composed of stars born around the solar circle.
@@
To summarise, the migrated sequence is composed of the age-metallicity relation at the initial corotation radius for the time before the bar formation and the segments of age-metallicity relations at the corresponding corotation radii at later times. The migrated sequence after the bar formation is flatter than the expected age-metallicity sequence at a fixed radius because stars at larger radii are dragged, but the extent of flattening depends on the details of the chemical evolution history (see Fig. A1 and Section 4.3.1 for clarification).
@@
In panel (c) of Fig. 1, we show the column-normalised age-metallicity distribution for the subgiant star sample described in Section 2. As for the test particle simulation, there are two distinct sequences marked by the blue and orange dashed lines. The test particle simulation results suggest that the lower sequence is composed of stars that formed around $R\sim R_{\odot}$, and the upper sequence is composed of stars formed in the inner Galaxy that have migrated with the expansion of the corotation radius. Xiang & Rix (2022) also pointed out the migrated origin of the upper sequence. Under this scenario, the turning point in the upper sequence, indicated by the red dashed line, gives an estimate of the bar formation epoch of $\sim 8$ Gyr ago, which is broadly consistent with other measurements of the bar formation time (Sanders et al., 2024; Haywood et al., 2024). We indicate the metallicity and age of the Sun with the red circle in panel (c). The Sun resides on the migrated sequence, suggesting that the Sun may have migrated to its current location with the corotation of the bar (Baba et al., 2024). A difference between panels (a) and (c) is that the upper envelope of the migrated sequence in the observations is decreasing whilst its counterpart in the test particle simulation is flat. This implies that either the deceleration rate is underestimated (see discussion in Section 4.3.1) or other non-axisymmetric perturbations neglected in the simulation have enhanced the radial migration.
@@ 4.2 Implication for the disc $[\alpha/\mathrm{Fe}]$-bimodality
Our results supplement previous analyses that have shown that radial migration can reproduce the $[\alpha/\mathrm{Fe}]$-bimodality in the Galactic disc (Sharma et al., 2021a; Chen et al., 2023). The scale length growth of the disc has to be more finely tuned to match the observed radial dependence of the $\mathrm{[\alpha/Fe]-[Fe/H]}$ plane for radial migration mechanisms that do not favour the current corotation radius. Our analysis suggests decelerating bar-driven radial migration is the key mechanism responsible for the disc $[\alpha/\mathrm{Fe}]$-bimodality, which causes the migrated and local stars to co-exist around the present corotation radius, naturally. However, other effects in addition to radial migration could also cause $[\alpha/\mathrm{Fe}]$-bimodality in the Galactic disc, and our results cannot rule out their existence (e.g. Khoperskov et al., 2021; Agertz et al., 2021; Renaud et al., 2021).
@@ 5.1 Predictions for the bar-driven radial migration scenario
Figure 5: Left: the fraction of outward migrated stars as a function of azimuthal angle in different guiding radii bins. The verticle solid and dashed lines denote the bar major and minor axes. Middle: the observed fraction of high-$\alpha$ stars as a function of azimuthal angle. Right: the radial surface density profile of the old stars/high-[Fe/H] stars in the test particle simulation with a decelerating bar (solid lines) and with a constant pattern speed bar (dashed lines). The verticle dashed line represents the current corotation radius.
@@
Azimuthal metallicity variation: Di Matteo et al. (2013) showed that ongoing radial migration can cause azimuthal abundance variations in the galactic disc. A key difference in the bar-driven radial migration scenario from other migration mechanisms is the azimuthal distribution of the migrated stars. Due to the characteristic shape of the corotation orbits, corotating stars naturally clump along the bar minor axis around the corotation radius. Therefore, the stars that experienced corotation dragging and migrated from the inner Galaxy would form an overdensity along the bar minor axis as shown in the left panel of Fig. 5, causing the mean metallicity to peak along the bar minor axis. To demonstrate this, we use our decelerating bar test particle simulation without the solar neighbourhood cut to show the fraction of outward migrated stars as a function of azimuthal angle in different guiding radii bins, as shown in the left panel of Fig. 5. The orange line illustrates a clear azimuthal dependence in the spatial distribution of the migrated stars around the corotation resonance. Azimuthal metallicity variation has been observed in the Milky Way disc (Poggio et al., 2022; Hawkins, 2023; Hackshaw et al., 2024). Recently, Khoperskov et al. (2024) measured the metallicity distribution of the entire Galactic disc with orbital reconstruction. As shown in fig. 22 therein, the metallicity for stars around guiding radii of $7-8$ kpc varies sinusoidally in the azimuthal direction and peaks along the bar minor axes. Such an azimuthal trend is not seen in stars at other guiding radii. This result is in agreement with the prediction of our model. However, kinematic fragmentation (Fragkoudi et al., 2017; Debattista et al., 2017; Zhang et al., 2024c) could also cause a similar result, as kinematically cold (more metal-rich) stars are easier to be trapped. Therefore, more analysis is needed to differentiate these mechanisms.
@@
Azimuthal variation in the high-$\alpha$ star fraction: As we discussed in Section 4.2, high-$\alpha$ stars observed in the solar vicinity are members of the migrated star sequence. Therefore, similarly to the azimuthal variation of metallicity, we might also expect an azimuthal variation for the high-$\alpha$ star fraction ($N_{\rm high-\alpha}$/$N_{\rm total}$), where more high-$\alpha$ stars are expected along the bar minor axes. As shown in the middle panel of Fig. 5, we examine this using a subsample of APOGEE DR17, following the quality cuts we addressed in detail in Appendix D. We indeed see an azimuthal variation in the fraction of high-$\alpha$ stars around the corotation radius, while the variation is much smaller outside the corotation radius. The fraction is smallest around the bar’s major axis and increases towards the minor axis. The prediction is also in agreement with the results shown in Khoperskov et al. (2024), who found that the high-$\alpha$ sequence contributes more in the direction of the minor axes than the major axes.
@@
Coldness of the Galactic disc: Frankel et al. (2020) found that diffusion in the angular momentum of the Galactic disc stars is much greater than the scattering of radial actions. Recently, Hamilton et al. (2024) showed that it is difficult to transport stars with such coldness with transient spiral arms unless the lifetime and pitch angle for the spiral arms are fine-tuned (but also see the counterevidence in Sellwood & Binney, 2025). However, in the bar-driven radial migration scenario, the angular momentum of the corotating stars increases with the slowing of the bar, but the radial actions are invariant (see Sellwood & Binney 2002; Chiba et al. 2021 for theoretical explanations). Computationally, Khoperskov et al. (2020) showed that the orbits of stars transported with the corotation resonance can remain circular. So, the bar-driven radial migration scenario can alleviate the problem of the coldness of the migration, but more experiments or quantitative analysis are required to demonstrate which fraction of the slowing bar-driven migration is needed to explain the observation.
@@
Density profile of the disc: As argued, the Galactic high-$\alpha$ and high-$\mathrm{Fe}$ disc should also experience significant radial migration driven by the decelerating bar. Stars that experienced resonant dragging clump around the present-day corotation radius producing a maximum in the radial exponential surface density profile of the high-$\alpha$ and high-$\mathrm{Fe}$ (super-solar) disc leading to the broken-power law profiles as observed by (Yu et al., 2021; Lian et al., 2022b). We demonstrate the role of the decelerating bar using our simulations in the right panel of Fig. 5. We show the surface density profile of the simulation with a decelerating bar (in solid lines) and with a constant pattern speed bar (in dashed lines) for an old population ($\rm age>8$ Gyr, potentially high-$\alpha$ stars in the Milky Way) and a high-[Fe/H] population ($\rm[Fe/H]>0.2$). The surface density profiles of both populations deviate from the initial exponential profile around the current corotation resonance in the presence of a decelerating bar. As other mechanisms may also lead to the broken power law density profile, more studies are required to build the unique correlation between the disc’s density profiles and the bar-driven radial migration mechanism.
@@ 5.2 Caveats of the test particle simulation
Second, the initial conditions of stellar particles released to the test particles after bar formation are in disequilibrium with the potential in the bar region because the initial conditions are axisymmetric while the potential is barred. This leads to an unrealistic simulation result in the bar region, but stellar particles born outside the bar region are less affected by this issue as the non-axisymmetric component of the potential becomes weaker. As we mainly analyse the corotation resonance and stars around it, which is constantly outside the bar region in the simulation setup, this issue would not impact our conclusion much.
Third, we adopted a specific chemical model to assign the metallicity of the particles in the simulation. The detailed quantitative results presented in this work depend on the method in Lu et al. (2024b). However, the nature of our argument is that the decelerating bar-driven radial migration causes stars from very different birth radii to co-exist around the corotation radius, and the metallicity of the migrated sequence naturally flattens after the bar formation because the corotation resonance moves outward. Hence, switching to a different chemical model does not affect our argument as long as the radial metallicity gradient in the model is reasonable. Furthermore, our motivational calculation in Section 4.3.1 and the observation of the break in the upper migrated age-metallicity sequence are direct empirical indicators which lead to very similar conclusions as the more detailed birth radii calculation.
@@ 6 Conclusion
We provided a series of other predictions if the slowing down of the Galactic bar is the main driver responsible for the radial migration in the Milky Way disc. We expect that: (1) the mean metallicity of stars is higher along the bar’s minor axis around the corotation radius; (2) the fraction of high-$\alpha$ stars is the smallest along the bar’s major axis and increases towards the minor axis around the corotation radius, which is demonstrated in Fig. 5 using APOGEE survey; (3) the migrated stars experienced much more radial migration than radial heating; (4) the radial density profile of the high-$\alpha$ and high-Fe stars should be consistent with a broken power law with the break close to the corotation radius. As discussed in Section 5.1, many of these predictions are consistent with observations in the Milky Way.
@@ A.1 Time evolution of the galactic potential
where the initial pattern speed is $\Omega_{i}=80$ km s-1 kpc-1, final pattern speed $\Omega_{4}=34$ km s-1 kpc-1, $t_{0}\sim 4$ Gyr, $t_{1}\sim 5$ Gyr, $t_{2}\sim 5.5$ Gyr, and $t_{4}\sim 11.2$ Gyr. We also adjust the length of the bar scaled to its corotation radius so that the length of the bar matches its original value when the pattern speed is $39\,\mathrm{km\,s^{-1}\,kpc^{-1}}$ (the pattern speed value of the Portail et al. (2017) model). At the end of the simulation, the pattern speed is $34\,\mathrm{km\,s^{-1}\,kpc^{-1}}$.
@@ Appendix D Azimuthal variation for high-$\alpha$ star fraction
We classify stars as a high-$\alpha$ star if $[\alpha/\rm M]>0.18$ for $[{\rm M/H}]<-0.7$ and $[\alpha/\rm M]>0.18-0.01([{\rm M/H}]+0.7)$ for $[{\rm M/H}]\geq-0.7$. We then take the ratio of the high-$\alpha$ to the total stars at different azimuthal angles in each of the guiding radii bins as shown in the middle panel of Fig. 5. There is an obvious increase of the high-$\alpha$ star fraction for the azimuthal angles away from the bar-major axis around the present-day guiding radii ($76<R_{g}/{\rm kpc}<$ 7), while the variation is much more insignificant inside and outside the corotation. This is in agreement with the expected consequence of the radial migration driven by the decelerating bar (see more discussion in Section 5.1).